the concentration in the IV bag is 150 mg in 200 ML. if the IV begins on monday at 0800 hours and must infuse with a dosage rate of 1.5 mg/min then
Answer #1

To find iv finishing time, first find out flow rate.

Available dose 150 mg

Available volume 200ml

Disired dose 1.5 mg/min

flow rate= (disired dose/ available dose)×available volume

= ( 1.5/150)×200

= 0.01×200

= 2ml/min

Multiply with 60min to convert ml/min to ml/hr

So ml/hr = 2×60

= 120ml/min

Infusion beginning time 08.00hrs

Available solution in the bag 200 ml

Flow rate(ml/min) = 2ml/min

Time required to finish 200ml=200/2

= 100min=1 hr 40 min

As iv start at 08.00hrs.it will finish at 09.40hrs